### v1

# Feedback Delay Network Optimization

Abstract
A common bane of artificial reverberation algorithms is spectral coloration, typically manifesting as metallic ringing, leading to a degradation in the perceived sound quality. This paper presents an optimization framework where a differentiable feedback delay network is used to learn a set of parameters to reduce coloration iteratively. The parameters under optimization include the feedback matrix, as well as the input and output gains. The optimization objective is twofold: to maximize spectral flatness through a spectral loss while maintaining temporal density by penalizing sparseness in the parameter values. A favorable narrower distribution of modal excitation is achieved while maintaining the desired impulse response density. In a subjective assessment, the new method proves effective in reducing perceptual coloration of late reverberation. The proposed method achieves computational savings compared to the baseline while preserving its performance. The effectiveness of this work is demonstrated through two application scenarios where natural-sounding synthetic impulse responses are obtained via the introduction of attenuation filters and an optimizable scattering feedback matrix.
keywords:
Audio systems, gradient methods, optimization, psychoacoustics, reverberation
1 Introduction
Delay-based recursive structures are a common method for generating artificial reverberation [1]. The initial approach incorporating delay lines and feedback loops was introduced by Schroeder and Logan in the early 1960s [2]. Over time, this structure underwent various modifications, eventually evolving into the feedback delay network (FDN) [3], which has since the 1990s become a widely adopted algorithm for digital reverberation synthesis [1].
A commonly encountered issue in delay-based artificial reverberation is coloration, often manifesting as audible metallic ringing in the resulting sound [2]. This coloration proves detrimental to perceived sound quality, deviating the reverb from the ideal standard of a smooth and spectrally flat reverberation, particularly in the tail. In its early form, Schroeder and Logan [2] sought to achieve colorless reverberation by concatenating delay-line-based allpass filters. Jot and Chaigne [3] introduced a two-stage design process involving the creation of a lossless system followed by the introduction of delay-proportional attenuation. However, these efforts proved insufficient, necessitating additional considerations regarding the selection of parameters to effectively reduce or eliminate this coloration effect [4, 5, 2, 6].
Recently, the perceived coloration was attributed to the modal properties of synthesized reverbs, with a wide distribution of modal excitations correlating positively with increased coloration [7]. Therefore, modal decomposition of artificial reverberation architectures, such as FDNs [8], can aid in identifying and mitigating phenomena that are linked to perceptual artifacts. In FDNs, the values of the modal excitations depend on all system parameters [8].
In this paper, the objective is to optimize the gain parameters and feedback matrix of an FDN. The goal is to minimize the perceptual coloration observed in its impulse response (IR). We adopt a stochastic gradient descent approach, focusing specifically on the magnitude of the frequency response. This paper extends our previous work [9] in which part of the loss function was based on the IR. The current work also demonstrates two practical applications of the proposed approach in synthesizing natural-sounding IRs by including attenuation filters in the FDN structure and optimizing the scattering feedback matrix. As we now concentrate solely on the frequency domain, the computational requirements are streamlined. We attain control over temporal density by utilizing its correlation with the density of the feedback matrix entries. Evaluation is conducted by analyzing the modal excitation distribution. A perceptual evaluation against the FDN design before optimization shows that the proposed method successfully decreased perceived coloration.
The paper is structured as follows. In Section 2, we provide background information on FDNs, including their modal decomposition, and discuss coloration in the IR of FDNs. Section 3 introduces our proposed method for designing colorless FDNs, highlighting the improvements made compared to our previous work. Practical applications of this approach are demonstrated in Section 4. The results of both objective and perceptual evaluations are presented in Section 5 and Section 6, respectively. Finally, Section 7 offers concluding remarks.
2 Background
In the following, we provide background information on the FDN and its fundamental properties. This includes an explanation of its modal decomposition and a discussion of coloration in FDNs, which are central concepts in the proposed method.
2.1 Feedback delay network
Figure 1: FDN with a single input, a single output, and an $N$-by-$N$ feedback matrix A. The thick lines indicate $N$-channel signal paths. The parameters under optimization are highlighted in blue.
An FDN is a recursive system consisting of delay lines, a set of gains, and a scalar feedback matrix through which the delay outputs are coupled to the delay inputs. Fig. 1 presents an example of a single-input, single-output (SISO) FDN architecture. The transfer function of the FDN can be written as
$H(z)=\frac{Y(z)}{X(z)}=\bm{c}^{\top}\big{[}\bm{D_{m}}(z)^{-1}-\bm{A}\big{]}^{-1}\bm{b}+d,$ (1)
where $\bm{A}$ is the $N\times N$ feedback matrix when $N$ is the number of delay lines, $\bm{D_{m}}(z)$ is the $N\times N$ delay matrix, vectors $\bm{b}$ and $\bm{c}$ are $N\times 1$ column vectors of input and output gains, respectively, the scalar coefficient $d$ is the direct gain, and the operator $(\cdot)^{\top}$ denotes the transpose. The frequency-domain input and output signals are represented by $X(z)$ and $Y(z)$, respectively. The vector $\bm{m}=[m_{1},\dots,m_{N}]$ defines the lengths of delays in samples. Its corresponding delay matrix $\bm{D_{m}}(z)$ is created by taking a diagonal matrix with entries given by $[z^{-m_{1}},\dots,z^{-m_{N}}]$. The sum of the delays gives the order of the system, i.e., $\mathcal{{M}}=\sum_{i=1}^{N}m_{i}$ [4].
The poles $\lambda_{i}$ of the system in (1) are the roots of its generalized characteristic polynomial (GCP) $p(z)$, and are fully characterized by $\bm{m}$ and $\bm{A}$:
$p(z)=\text{det}(\bm{D_{m}}(z)^{-1}-\bm{A})\;.$ (2)
With (2), the transfer function (1) can be expressed as a rational polynomial
$H(z)=d+\frac{\bm{c}^{\top}\text{adj}(\bm{D_{m}}(z)^{-1}-\bm{A})\bm{b}}{p(z)},$ (3)
where $\text{adj}(\cdot)$ indicates the matrix adjugate operator. This expression of the FDN transfer function can be used to compute the modal decomposition of the system directly. Moreover, the GCP provides explicit information regarding the stability of the system, which will subsequently be utilized for parameter-tuning purposes.
2.2 Modal decomposition
The IR of the FDN can be expressed as the sum of complex one-pole modes, also known as resonators [8]:
$h(n)=\sum_{i=1}^{\mathcal{{M}}}h_{i}(n)\;.$ (4)
Each mode $h_{i}(n)$ is defined by the pole $\lambda_{i}$ and the residue $\rho_{i}$:
$h_{i}(n)=|\rho_{i}||\lambda_{i}|^{n}e^{\jmath(n\angle\lambda_{i}+\angle\rho_{i})}\,,$ (5)
where ||$\cdot$|| ⋅ | denotes the magnitude, $\angle$ represents the argument of a complex number in radians, $\jmath=\sqrt{-1}$, and $n$ indicates the discrete time index. The total number of poles $\mathcal{{M}}$ is the same as the system order.
The transfer function of the FDN (1) can be represented by its poles and residues through the partial fraction decomposition as
$H(z)=d+\sum_{i=1}^{\mathcal{{M}}}\frac{\rho_{i}}{1-\lambda_{i}z^{-1}}\;,$ (6)
commonly referred to as the modal decomposition of the FDN [8]. The excitation and initial phase of the $i^{\text{th}}$ mode are determined by the magnitude $|\rho_{i}|$ and phase $\angle\rho_{i}$, respectively, of its corresponding residue. The magnitude and phase of the $i^{\text{th}}$ pole, $|\lambda_{i}|$ and $\angle\lambda_{i}$, respectively, dictate its decay rate and frequency.
A common approach to calculate the modal decomposition (6) from the rational polynomial form of the transfer function (3) involves determining the system poles $\lambda_{i}$. These poles are obtained as the roots of the GCP (2), derived from the eigenvalues of the linearized feedback matrix [8].
2.3 Homogeneous decay in FDN
Designing an artificial reverberator with FDN often starts by creating a lossless prototype, having an energy-preserving feedback loop [10, 11]. The FDN is said to be lossless if the roots of the GCP (2) have magnitude equal to one, i.e., $|\lambda_{i}|=1$ for all $i$ [12]. The advantage of initially designing a lossless FDN lies in the straightforward implementation of frequency-dependent decay that equally influences all system poles. This is achieved by extending every delay with a frequency-dependent attenuation filter to meet the specified reverberation time [3].
In this study, we mainly focus on the specific case of frequency-independent homogeneous decay. This refers to the case where all modes experience the same decay rate, i.e., $|\lambda_{i}|=\gamma$ for all $i$’s. In Subsection 4.1, we will demonstrate how an FDN optimized at such $\gamma$ can be employed as a lossless prototype for synthesizing reverb with frequency-dependent decay.
Homogeneous decay is achieved with a feedback matrix $\bm{A}$ being the product of a unilossless matrix $\bm{U}$ and a diagonal matrix $\bm{\Gamma}$, whose entries are delay-proportional absorption coefficients, $\bm{\Gamma}=\text{diag}(\gamma^{\bm{m}})$. The feedback matrix can be expressed as
$\bm{A}=\bm{U}\bm{\Gamma}\;.$ (7)
A matrix $\bm{U}$ is unilossless if, regardless of the choice of delays $\bm{m}$, its eigenvalues are unimodular and its eigenvectors are linearly independent. A matrix $\bm{U}$ satisfying the unitary condition, $\bm{U}\bm{U}^{H}=\bm{I}$, is also unilossless [6, 13]. As $\bm{U}$ is unilossless, the modal decay is controlled entirely by the gain-per-sample parameter $\gamma$, where $0\leq\gamma\leq 1$. The gain-per-sample in dB is
$\gamma_{\mathrm{dB}}=\frac{-60}{f_{\textrm{s}}T_{60}}\;,$ (8)
where $f_{\textrm{s}}$ is the sampling rate in Hz and $T_{60}$ is the reverberation time in s defined as the time required for the sound level to decay by $60$ dB from the initial steady-state value.
2.4 Coloration in FDN
Figure 2: Magnitude response of an FDN with size $N=8$, showcasing the difference between long and short delay lines when the former is 10 times longer than the latter.
Natural late reverberation can be conceptualized as white noise with an exponential decay, providing perceptually ideal smooth diffuse reverberation [14]. In artificial reverberation, achieving a white noise generator involves ensuring the allpass property is maintained. Various perceptual artifacts of recursive reverberators can be attributed, though not exclusively, to deviations from this ideal, which then result in unwanted coloration.
Schroeder and Logan [2] made the initial attempt to produce colorless artificial reverberation by establishing specific requirements for the reverberators in addition to a flat frequency response. Overlapping normal modes across all frequencies, equal $T_{60}$ values for each mode, sufficient echo density, lack of periodicity in the time domain, and no periodic or comb-like frequency responses were deemed necessary to achieve colorlessness [2]. Despite fulfilling the aforementioned conditions, however, the Schroeder series allpass did not attain complete colorlessness.
A recent study was conducted to understand further the role of modal excitation in late reverberation coloration [7]. Listening test results suggest that a narrow distribution of the modal excitation values $|\rho_{i}|$ tends to result in a perceptually white spectrum [7]. For large values of $|\rho_{i}|$, coloration starts to become noticeable [7]. We also observed this outcome in our earlier study, where we conducted a listening test focusing on the coloration of FDN-based reverb [9]. In this test, we compared the reverb before and after applying an optimization method to narrow the modal excitation distribution.
The literature also indicates that more than 6000 modes are needed for an IR to be perceived as rather colorless [15]. Additionally, it was shown that for large values of $\mathcal{{M}}$, the modes of the FDNs are uniformly distributed [8], preventing additional coloration that usually results from clusters of modes. Nonetheless, a flat magnitude response and a uniform modal frequency distribution are insufficient to achieve colorlessness.
To comprehend the complexities associated with coloration in FDN, we can examine the comb-filter structure as a special case. Specifically, when the feedback matrix $\bm{A}$ is diagonal, the FDN adopts the configuration of a parallel comb-filter structure. In the case of a homogeneous FDN with a diagonal feedback matrix, the transfer function in (1) is analogous to a superposition of comb filters, with each filter having the transfer function
$H_{i}(z)=\frac{1}{1+\gamma^{m_{i}}z^{-m_{i}}}\,.$ (9)
The contribution of each filter to the total energy of the response can be calculated as
$\lVert H_{i}(z)\rVert_{2}^{2}$ $=\int_{0}^{2\pi}\lvert H_{i}(e^{\jmath\omega})\rvert^{2}\textrm{d}\omega$ (10)
$=\frac{1}{1-\gamma^{2m_{i}}}\,,$ (11)
where $\omega$ is the angular frequency, and $\lVert\cdot\rVert_{2}$ denotes the $\ell^{2}$ norm.
For any given value of $\gamma$, (11) represents an exponential function. For positive $m_{i}$, the function approaches infinity as $m_{i}$ values decrease and tends to zero as $m_{i}$ values increase. For example, let us consider the case when the gain-per-sample is $\gamma=0.9999$ ($T_{60}=1.44$ s). In this scenario, the energy of the response from the comb filter with $m_{1}=200$ is 8.5 times greater than that with $m_{2}=10m_{1}=2000$. Specifically, $\lVert H_{1}(z)\rVert_{2}^{2}=25.5$, while $\lVert H_{2}(z)\rVert_{2}^{2}=3$. Fundamentally, shorter delays contribute more energy and produce strongly audible metallic-sounding comb peaks. In contrast, longer delays contribute less energy and tend to be masked by the more dominant comb filters [16].
Figure 2 shows a section of the magnitude spectrum of two FDNs with identical parameters but different delay-line lengths. The FDN with shorter delay lengths, approximately one-tenth of the longer delays, exhibits higher magnitude values and more distinct, louder resonances. For long delays, the resonances become less noticeable, and the loudest frequencies exhibit a magnitude approximately 20 dB lower compared to the scenario with shorter delays. This example shows that to achieve a colorless FDN, it is necessary to avoid strongly recirculating short delays and encourage exciting long delays instead.
2.5 Problem statement
In this paper, our goal is to optimize the feedback delay matrix $\bm{A}$, as well as the input and output gains $\bm{b}$ and $\bm{c}$, to enhance the perceptual colorlessness of the resulting IR. In Fig. 1, the FDN parameters being optimized are indicated in blue. Previous studies indicate that the choice of frequency-dependent attenuation has little impact on coloration [7]. Therefore, we conduct optimization on a long-ringing, frequency-independent prototype FDN. For ideal late reverberator synthesis with this setup, the spectrum’s magnitude must remain constant across frequencies and time. However, the recursive structure of FDN introduces undesired colorations, which constitutes a primary challenge in our work, as emphasized in Subsection 2.4.
Another challenge is maintaining the temporal density of the IR during optimization. Our previous work demonstrated that focusing solely on the frequency-sampled magnitude response may lead to convergence towards a comb filter [9]. To address this efficiently, we utilize the relationship between temporal density and the density of the feedback matrix. For evaluation, we adopt modal decomposition of the FDN transfer function, as it has proven useful in analyzing the coloration of recursive systems [7]. To demonstrate the effectiveness of our method, we present various application scenarios that showcase its utilization for frequency-dependent attenuation and increased echo density, all while ensuring that the computational efficiency is maintained.
3 FDN optimization
In this section, we present a method to reduce coloration in the IR of an FDN with frequency-independent $T_{60}$. We employ stochastic gradient descent to optimize the parameters of the FDN, which has been made differentiable by utilizing the frequency sampling method.
3.1 Differentiable FDN
Using the frequency sampling method, the FDN is approximated as an FIR filter. An FIR filter is straightforward to implement, as it does not require a recursive structure. Moreover, an FIR filter can be easily parallelized and optimized for hardware acceleration, such as a graphics processing unit (GPU). To obtain the FIR approximation of $H(z)$, the delay matrix $\bm{D_{m}}(z)$ is evaluated at discrete frequency points
[…]
where $M$ indicates the total number of frequency bins equally distributed on the unit circle.
The block diagram of the architecture is depicted in Fig. 3. We integrated $H(\bm{z}_{M})$ into an optimization framework to estimate a set of FDN configurations based on a spectral loss and a sparsity loss by gradient descent. The learnable parameters are the feedback matrix $\bm{A}$ and the input and output gain vectors $\bm{b}$ and $\bm{c}$, respectively. The delay lengths $\bm{m}$ are set at initialization and kept constant during training. The direct gain $d$ is set to zero. During training, the FDN is set to have a homogeneous decay by forcing $\bm{A}$ to satisfy (7) for a given $\gamma$.
Figure 3: Architecture of the proposed optimization workflow. A thick line indicates a $N$-channel signal path. Backpropagation is highlighted in blue.
At each training step the estimated channel-wise transfer function $\bm{H}(z)$ is computed at a batch of frequency bins,
$\bm{H}(\bm{z}_{\mu})=\bm{c}\big{[}\bm{D_{m}}(\bm{z}_{\mu})^{-1}-\bm{A}\big{]}^{-1}\bm{b}+d\,.$ (13)
Every batch is composed of $\mu<M$ different frequency points that are randomly sampled from $\bm{z}_{M}$. The value of $M$ is such that to ensure oversampling, while we found empirically that $\mu$ can be as small as $0.05f_{\textrm{s}}$. At each epoch, $\bm{H}(z)$ is estimated at all frequency points within the vector $\bm{z}_{M}$. In our prior study [9], the computation of $\bm{H}(z)$ during each training step involved the entire vector $\bm{z}_{M}$ rather than a subset, resulting in a bottleneck due to the computational complexity associated with matrix inversion and multiplication.
3.2 Orthogonal feedback matrix
To guarantee frequency independence of the energy decay, we adopted the parameterization in (7), where the feedback matrix is computed from an orthogonal matrix $\bm{U}$. The class of orthogonal matrices meets the unitary condition for losslessness and can be obtained through a parameterization of skew-symmetric matrices.
At initialization, a random matrix $\bm{W}$ is constructed. At each optimization step, $\bm{W}$ is mapped to a skew-symmetric matrix, and the matrix exponential is computed,
$\bm{U}=e^{\bm{W}_{\textrm{Tr}}-\bm{W}_{\textrm{Tr}}^{\top}}\,,$ (14)
where $\bm{W_{\textrm{Tr}}}$ is the upper triangular part of $\bm{W}$ and the operator $e^{(\cdot)}$ denotes the matrix exponential. The mapping in (14) implicitly ensures orthogonality of $\bm{U}$ and can be used in regular gradient descent optimizers without creating spurious minima [17]. During training, the backward pass differentiates through the parameterization in (14) and $\bm{W}$ is updated accordingly.
3.3 Gain-per-sample
Figure 4: Magnitude response of a homogeneous FDN of order $\mathcal{M}=8944$ with random orthogonal feedback matrix and unitary input/output gains at different values of $T_{60}$ in seconds. In the lossless case, $T_{60}=\infty$, the resonances tend to infinity.
For stable training, the feedback matrix $\bm{A}$ must have eigenvalues with a modulus less than one, indicating losses. In the lossless case, i.e., $\lvert\lambda_{i}\rvert=1$, evaluating $H(\bm{z}_{M})$ becomes infeasible, as the denominator in (3) becomes null.
In our work, we adopted the homogeneous FDN where $\bm{A}$ is parameterized according to (7), and $\gamma$ is set at initialization to a value lower than one which is constant during optimization. The value of $\gamma$ used during optimization is chosen by examining the connection between the mean damping factor $\overline{\delta}$, used in room acoustics, and the mean spacing of resonance frequencies $\overline{\Delta f}$ To guarantee that the modes are well separated, the mean spacing of resonance frequencies should be larger than the average resonance half-width [18]
$\overline{\Delta f}\gg\frac{\overline{\delta}}{\pi}\;.$ (15)
In room acoustics, the limiting frequency below which the modes are well-separated is called Schroeder frequency, indicated here as $f_{\textrm{Schroeder}}$ [19]. This frequency marks the threshold above which an average of at least three modes falls within one resonance half-width. Using the fact that in FDNs the modal frequencies are nearly equally distributed [8], we can derive the limiting average resonance half-width
$\overline{\Delta f}_{\mid f=f_{\textrm{Schroeder}}}=3\frac{f_{\textrm{s}}}{\mathcal{{M}}}\;.$ (16)
We can use the above conditions to determine the minimum value for $T_{60}$ to be used during training
$T_{60}\gg\frac{\mathcal{{M}}\text{ln}(10)}{\pi f_{\textrm{s}}}\;.$ (17)
Increasing the value of $T_{60}$ leads to modes with lower half-widths and greater separation between them. For a target $T_{60}$, the value of $\gamma$ can be derived from (8). However, as $\gamma$ approaches 1, the resonance peaks in the magnitude response become narrow, making it impossible to obtain a flat magnitude response by combining the resonances.
Figure 4 shows the effect of increasing $\gamma$ on the resonance width in a short section of the FDN magnitude response. To enhance the interpretability of the figure, its legend reports the $T_{60}$ values associated with $\gamma$. The sharp peaks visible when $T_{60}=\infty$ ($\gamma=1$) are significantly smoothed in the other two curves ($\gamma<1$). For $T_{60}=0.14\,$s ($\gamma=0.999$), the resonance half-widths become excessively large, making it difficult to identify individual resonances.
We ran optimization at different values of $\gamma$, and we observed good convergence when $T_{60}\leq 10$ s. At inference, $\gamma$ serves as a free parameter, enabling the generation of reverberation with arbitrary $T_{60}$ values.
3.4 Parameter initialization
In stochastic gradient descent, as in many other optimization algorithms, the choice of initial parameter values can affect the algorithm’s ability to find the global or a good local minimum of the loss function. To start from a point that facilitates an efficient exploration of the loss landscape, we initialize the values of $\bm{b}$ and $\bm{c}$ by drawing them from the normal distribution $\mathcal{N}(0,N^{-1})$, and the entries of $\bm{W}$ from the uniform distribution $\mathcal{U}(-1/\sqrt{N},1/\sqrt{N})$.
The design of the delays is a rather non-trivial task that requires further constraints. To maximize the echo density, the delay lengths should be co-prime [16]. However, the concentration of delays around a certain value may lead to a perceivable strong energy fluctuation over time. Moreover, low-order dependencies, which are integer linear combinations of delays that coincide with other integer linear combinations of delays with small coefficients, can also contribute negatively to the smoothness of the response [6]. To avoid degenerative patterns and ensure a smooth-sounding reverb, we choose delays that are logarithmically distributed co-prime numbers leading to $\mathcal{{M}}\geq 6000$.
3.5 Loss function
Figure 5: Evolution of losses (left) and the relative density increase of the IR (right).
The network is trained on two losses, $\mathcal{L}_{\textrm{spectral}}$ and $\mathcal{L}_{\textrm{sparsity}}$. The spectral loss $\mathcal{L}_{\textrm{spectral}}$ aims to minimize the frequency-domain mean-squared error between the absolute value of the predicted magnitude response for each channel and the target flat magnitude response. The sparsity loss $\mathcal{L}_{\textrm{sparsity}}$ penalizes sparseness in the parameter values to encourage density in the time domain. The total loss function is
$\mathcal{L}=\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))+\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ (18)
and the individual loss terms are
$\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))$ $=\frac{1}{\mu}\sum_{z\in\bm{z}_{\mu}}\sum_{i=1}^{N}\left({\lvert H_{i}(z)\rvert-1}\right)^{2}$
$+\left({\lvert H(z)\rvert-1}\right)^{2}$
$\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ $=\frac{N\sqrt{N}-\sum_{i,j}\lvert U_{ij}\rvert}{N(\sqrt{N}-1)}$
where $H_{i}(z)$ is the output of the network’s $i^{\text{th}}$ channel computed from the output of the $i^{\text{th}}$ delay line and scaled by $c_{i}$.
The sparsity loss, which is the last term above, is calculated from the sum of the absolute values of the matrix entries, i.e. $\sum_{i,j}\lvert U_{ij}\rvert$. For orthogonal matrices, its value is bounded by $-N\sqrt{N}$ and $N$. To make the dynamic range of the $\mathcal{L}_{\textrm{sparsity}}$ comparable with that of $\mathcal{L}_{\textrm{spectral}}$ we shifted $\sum_{i,j}\lvert U_{ij}\rvert$ by $N$ and normalized it by $N(\sqrt{N}-1)$. The term is then shifted by $1$ to span positive values only.
In our previous work [9], the sparseness was controlled by a temporal loss consisting in the ratio of the $\ell_{1}$ norm to the $\ell_{2}$ norm of the estimated IR. This required the inverse Fourier transform, slowing down the optimization steps.
To visualize the impact the loss function (18) has on the IR density, Fig. 5 illustrates how the $\ell_{1}$ norm of the estimated IR signal $h$, i.e., $\sum_{k}\lvert h(k)\rvert$, evolves in relation to the losses at each epoch of one run of training. The right pane shows the increase in density relative to initialization. The left pane shows the evolution of the loss and its terms at each epoch. The terms decay at similar rates, and after 10 epochs, they start to converge. Compared to the losses, the increase in density starts to converge quicker, reaching a $16\%$ increase after just 5 epochs. Without sparsity control, the magnitude response may converge towards that of a comb filter, leading to a corresponding decrease in IR density, as demonstrated in our previous work [9].
4 Applications
This section showcases how the proposed approach can be effectively employed in practical applications. Specifically, the differentiable FDN (DiffFDN) can be utilized for synthesizing reverberation with a frequency-dependent decay rate and creating dense reverb through the use of attenuation filters and an optimizable scattering feedback matrix.
4.1 Frequency-dependent decay control
To generate a natural-sounding room IR that closely matches a measured reference IR, attenuation filters are added to a lossless FDN to control the reverberation time as a function of frequency $T_{60}(\omega)$ [20, 21, 11, 22]. Using a differentiable FDN optimized for colorlessness as a lossless prototype can address coloration in the prototype that cannot be effectively mitigated by the attenuation filter, facilitating a reduction in size as well.
The attenuation filter is designed to approximate a target frequency-dependent reverberation time $T_{60}(\omega)$ by achieving a target gain-per-sample $\gamma(\omega)$
$\gamma_{\textrm{dB}}(\omega)=-\frac{60}{f_{\textrm{s}}T_{60}(\omega)}\,.$ (19)
Attenuation filters are commonly introduced after the delay lines [20]. To simplify the transfer function, at the price of a negligibly lower $T_{60}(\omega)$ accuracy, we place the filters in the feedback loop, resulting in the following transfer function
$H(z)=\bm{c}^{\top}\big{[}\bm{D_{m}}(z)^{-1}-\bm{U}\bm{\Gamma}(z)\big{]}^{-1}\bm{b}+d,$ (20)
where $\bm{\Gamma}(z)$ is the diagonal attenuation matrix, whose diagonal entries are the delay line specific per-sample attenuation filters $\Gamma_{i}(z)$. Ideally, $\bm{\Gamma}(z)$ would be zero-phase, such that the decay rate of the FDN is determined by $\lvert\bm{\Gamma}(z)\rvert^{1/\bm{m}}$ [3]. In practical designs, the attenuation filter is not zero-phase, introducing a group delay that is, however, small compared to the delays $\bm{m}$, hence negligible [11, 3].
We construct the attenuation filters as a graphic equalizer composed of a cascade of eight peaking filters and of a combination of second-order low-shelving and high-shelving filters for controlling the gains at DC and Nyquist limit, respectively [23]. The peaking filters have center frequencies at one-octave intervals ranging from 63 Hz to 8000 Hz. The low-shelving filter’s crossover frequency is set at 46 Hz, and the high-shelving filter’s crossover frequency is set at 11360 Hz.
We used DecayFitNet, a neural-network-based approach [24], to estimate each band’s $T_{60}$ and decay amplitude $A$ from a reference room IR. DecayFitNet estimates the parameters from the energy decay function (EDF) modeled as a sum of at most three multiple exponential decays and one noise term. In this work, we assume single slope decay, and we neglect the noise term. Each delay line’s output has a per-sample attenuation filter in the feedback path. This filter is designed to compensate for the delay $m_{i}$ introduced by the delay line by adjusting its target magnitude response
$\lvert\Gamma_{i}(e^{\jmath\omega})\rvert=10^{m_{i}\gamma_{\textrm{dB}}(\omega)/20}\,.$ (21)
The attenuation filters modify the energy of each mode. To align it with the target frequency response envelope, we design a shared graphic equalizer that matches the initial amplitude provided by the DecayFitNet and is placed after the output gains.
4.2 Optimizing scattering feedback matrix
A central challenge in the design of FDNs is the generation of sufficient echo density in the IR while maintaining computational efficiency [25]. To make the echo density grow faster in time and reproduce a scattering-like effect, the feedback matrix can be generalized to a filter feedback matrix (FFM) [26] where each entry is an FIR filter. The transfer function of the FDN becomes
$H(z)=\bm{c}^{\top}\big{[}\bm{D_{m}}(z)^{-1}-\bm{A}(z)\big{]}^{-1}\bm{b}+d,$ (22)
where $\bm{A}(z)=\bm{U}(z)\bm{\Gamma}(z)$, and $\bm{U}(z)$ is a FFM, which can be expressed in terms of scalar coefficient matrices $\bm{U}_{0},\bm{U}_{1},\dots,\bm{U}_{S}$, i.e.,
$\bm{U}(z)=\bm{U}_{0}+z^{-1}\bm{U}_{1}+z^{-2}\bm{U}_{2}+\dots z^{-S}\bm{U}_{S}.$ (23)
where $S$ is the maximum filter order of $\bm{U}(z)$ if $\bm{U}_{S}\neq 0$. The losslessness condition is satisfied if $\bm{U}(z)$ is paraunitary matrix, i.e., $\bm{U}(z^{-1})^{H}\bm{U}(z)=\bm{I}$, where $\bm{I}$ is the identity matrix and $(\cdot)^{H}$ denotes the complex conjugate transpose [12]. In this work, $\bm{U}(z)$ is realized as a paraunitary FIR filter using the following factorization:
$\bm{U}(z)=\bm{D}_{\bm{m}_{K}}(z)\bm{U}_{K}\cdots\bm{U}_{2}\bm{D}_{\bm{m}_{1}}(z)\bm{U}_{1}\bm{D}_{\bm{m}_{0}}(z),$ (24)
where $\bm{U}_{1},\dots,\bm{U}_{K}$ are $N\times N$ unitary matrices and $\bm{m}_{0},\dots,\bm{m}_{K}$ are vectors of $N$ integer delays [26]. In this arrangement, the FFM primarily incorporates $K$ delays and mixing stages into the main FDN loop.
Similar to the scalar feedback matrix, we optimize the unitary matrices and input and output gains to reduce the coloration. We construct the unitary matrices $\bm{U}_{k}$ as the coefficients of a $K\times N\times N$ tensor $\bm{W}_{k}$ parametrized according to (14) to ensure orthogonality. The delays $\bm{m}_{k}$ are fixed, and their values are selected to be relatively small in comparison to the primary delays $\bm{m}$ to add short-term density.
4.3 Optimizing Householder matrix
When using the Householder matrix $\bm{U}=\bm{I}-2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}$, where $\bm{v}$ is a unit vector, gives a simple, yet restricted, parametrization of orthogonal matrices. Using the Sherman-Morrison formula, i.e.,
(${\lparen\bm{W}+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}\rparen}^{\text{-}1}={\bm{W}}^{\text{-}1}-\frac{2{\bm{W}}^{\text{-}1}\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{W}}^{\text{-}1}}{1+2\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{W}}^{\text{-}1}\bm{v}}$ (25)
can help to avoid the matrix inverse. Thus, the FDN transfer function is then written as
$H(z)=\bm{c}^{\top}{\left({\bm{W}(z)+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}}\right)}^{\text{-}1}\bm{b}\,,$ (26)
where $\bm{W}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{I}$. With this parameterization, the computational complexity required to calculate the FDN transfer function at one frequency point is reduced from $\mathcal{O}(N^{3})$ to $\mathcal{O}(N)$.
To ensure homogeneous absorption, $\bm{\Gamma}$ is included in (26) as
$H(z)=\bm{c}^{\top}{\left({\bm{W_{\gamma}}(z)+2\bm{v}\bm{v_{\gamma}^{\mathstrut\scriptscriptstyle\top}}}\right)}^{\text{-}1}\bm{b}\,,$ (27)
where $\bm{W_{\gamma}}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{\Gamma}$ and $\bm{v_{\gamma}}^{\mathstrut\scriptscriptstyle\top}=\bm{v}^{\mathstrut\scriptscriptstyle\top}\Gamma$. With this parametrization, the optimizable parameters become $\bm{v},\,\bm{b},\,\text{and}\,\bm{c}$. The input to the sparsity loss remains $\bm{U}$, which is now defined using the Householder transformation.
5 Objective evaluation
The following section presents the FDN configuration and the objective evaluation of the proposed method. The assessment examines the modal excitation distribution.
5.1 Evaluation setup
We concentrate on smaller FDNs with limited delay lines, as these are significantly impacted by coloration. Specifically, we consider FDN sizes of $N=4,6,8$ The values of the delay-line lengths are presented in Table 1. The delay lengths are prime numbers distributed logarithmically. In all configurations, the total number of modes is $90008700<\mathcal{M}<$ 9000.
Table 1: Delay-line lengths for each size $N$ of the analyzed FDNs. Values are logarithmically distributed prime numbers.
N Lengths
4 [1499, 1889, 2381, 2999]
6 [997, 1153, 1327, 1559, 1801, 2099]
8 [809, 877, 937, 1049, 1151, 1249, 1373, 1499]
[…]
During optimization, we used a sampling rate of $f_{\textrm{s}}=48$ kHz and $M=480000$ frequency points equally distributed over $[0,\pi]$. The batch size was $\mu=2000$, requiring 240 steps to complete an epoch. For training, we used $80\%$ of the frequency points, while the remaining $20\%$ were used for validation. We employed Adam optimizer [27] with learning rate of $\eta=10^{-3}$. Regarding the gain-per-sample, we found that $\gamma=0.9999$ leads to stable training and fast convergence. This choice implies $T_{\textrm{60}}=1.439\,$s and satisfies (17).
The configurations under analysis include the differentiable FDN optimized according to our prior work (DAFx23) [9], the one optimized with the proposed method (DiffFDN), and two variations of the latter: one with an optimizable Householder matrix (HH) and another with an optimizable scattering matrix (SCAT).
5.2 Computational complexity
The presented optimization utilizes a sparse set of frequency sampling points at each step. This choice is driven by the aim to decrease the overall number of operations, providing a notable reduction when compared to the baseline method [9]. During the forward pass, the computational complexity needed to calculate the FDN transfer function at a single frequency point in (13) is $\mathcal{O}(N^{3})$. For each epoch, the proposed method (DiffFDN) computes the transfer function at $M$ points. In contrast, the baseline approach (DAFx23) requires a similar number of frequency points at each training step, resulting in a linear increase in the number of operations with the dataset size, e.g., 256 in [9]. This is because the number of epochs required for convergence in the baseline is close to that of the presented method. Additionally, in the forward pass of the baseline, the FFT is necessary due to the temporal loss, introducing additional operations. By employing the Householder matrix, we can significantly decrease the computational complexity of (13) to $\mathcal{O}(N)$. Introducing the FFT would raise this complexity to $\mathcal{O}(N\text{log}_{2}(N))$.
The FDNs also have different computational complexity during operation. For an FDN with $N$ delay lines, each equipped with an octave graphic equalizer, the number of multiply-and-add operations per sample is as follows: $2N$ operations for input and output gains $\bm{b}$ and $\bm{c}$, $2N$ for operations the delay lines, $44N$ operations for the attenuation filters, $N^{2}$ operations for the standard matrix multiplication, and $2N$ for the Householder matrix multiplication. The scattering matrix multiplication with $K$ stages are then $K(N^{2}+2N)$ operations.
Table 2: Number of operations for an FDN during operation at different FDN sizes and matrix types. In comparison, a typical standard FDN with size $N=32$ costs 2560 operations.
FDN type 4 6 8
RO 208 324 448
DiffFDN 208 324 448
DiffFDN (HH) 200 300 400
DiffFDN (SCAT) 288 480 704
5.3 Modal excitation distribution
To evaluate the improvements in coloration, we analyze the modal excitation distribution before and after the optimization. To obtain the residues $\rho_{i}$, we compute modal decomposition in (6) using the Ehrlich-Aberth iteration method [28, 8]. The method enables solving the eigenvalue problem for large delays $\bm{m}$ by finding an approximation to the eigenvalues based on the polynomial matrix formulation of the FDN [28].
Table 3 shows the standard deviation of the modal excitation distribution in dB, $20\text{log}_{10}\lvert\rho_{i}\rvert$, for each tested configuration. The numbers refer to 100 iterations, each with a unique set of randomly sampled initial parameters. Each optimization run is 20 epochs long, as our experiments showed no further improvement with extended training. In each scenario, the standard deviation decreases, signifying a reduction in coloration. Specifically, FDNs with a size of $N=4$ exhibit a more noticeable distribution narrowing. Based on this test alone, it is challenging to predict the overall trend across different FDN sizes.
Table 3: Standard deviation of the modal excitation before (Init) and after optimization (Optim) at different FDN sizes.
N Init Optim
4 7.8346 4.4518
6 7.8570 5.4239
8 8.0322 5.7813
[…]
To ensure that the changes to the optimization framework did not reduce the performance of the method introduced in [9], we conducted identical tests using our previous optimization framework. Each set of parameters used at initialization is shared between the two methods. We observed comparable convergence rates for both optimization methods. Consequently, we allowed each training run to complete 20 epochs before halting it and extracting the output parameters for evaluation.
The standard deviation of the model excitation before and after optimization is presented in Table 4. For these evaluations, we utilized an FDN of size $N=4$ with delays specified in Table 1. The proposed method shows results comparable to our previous work (DAFx23). Even though it is likely impossible to perceive a difference of 0.03 dB in standard deviation, the more significant improvement lies in the reduction of operation count.
Table 4 also shows the results of a similar test run on the FDN with optimizable scattering matrix (SCAT) and Householder matrix (HH), both of size $N=4$. Due to a different parameterization of the feedback matrix, the distributions of the initial parameters are different from the standard configuration. After optimization, the standard deviation reduces in both cases. Among all tested configurations, the HH case exhibited the smallest improvement, likely due to its more restrictive parameterization. The SCAT configuration also showed a lower degree of improvement compared to the standard configuration. This could be attributed to the higher complexity of the problem, suggesting potential enhancement through fine-tuning the optimization framework.
Table 4: Standard deviation of the modal excitation before (Init) and after optimization (Optim) of different frameworks, each employing an FDN of size N=4 and identical set of dealys.
FDN type Init Optim
DAFx23 7.8346 4.4802
DiffFDN 7.8346 4.4518
DiffFDN (HH) 8.5641 7.8765
DiffFDN (SCAT) 7.6200 5.8822
[…]
6 Perceptual evaluation
To further validate our results, a formal listening test was conducted to evaluate the perceived coloration in the IRs of the optimized FDN and quantify the improvement from IRs produced with the initial parameters.
6.1 Listening test procedure
Figure 6: Listening test results on the analysized reverberation conditions: (a) lossless, (b) homogeneous decay, and (c) frequency-dependent decay.
The test followed the Multiple Stimuli with Hidden Reference and Anchor (MUSHRA) recommendation [29], and it was carried out using the web audio API-based experiment software webMUSHRA developed by International Audio Laboratories Erlangen [30].
On each page, the listening test compared five sets of FDN configurations against a reference. Similarly to the objective evaluation, the test items included three configurations, i.e., three FDN sizes ($N=$ 4, 6, and 8). The configurations were evaluated on three different reverberation conditions, each covering a dedicated section of the test. The tested reverberation conditions are: lossless, homogeneous decay, and frequency-dependent decay. At the beginning of each part, a training page was presented to familiarize the subjects with the sound samples. Adjustments to the overall loudness were allowed during the initial training page but were maintained constant throughout the remainder of the test.
During the evaluation, participants rated the similarity between each presented item and the reference sound using a scale ranging from 0 to 100. On each page, six IRs were assessed, including an anchor and the hidden reference. To encourage subjects to assess samples based on coloration rather than subtle temporal features, the hidden reference differed from the actual reference as a distinct instance of the generated signal while maintaining its statistical properties.
In the part testing the lossless condition, we evaluated late non-decaying reverb. The reference for this part was a 3-second segment of white Gaussian noise, known to be highly colorless. The IRs of the FDNs were extracted from the mixing time, thereby eliminating the initial echo buildup from the evaluation. The section testing homogeneous decay, focused on frequency-independenty decaying IRs, with a $T_{60}$ value set at 2 seconds for all frequencies. A corresponding exponential decaying envelope with the same $T_{60}$ value was applied to the reference white Gaussian noise. The part testing frequency-dependent decay involved decaying IRs with a fixed frequency dependent $T_{60}(\omega)$ across conditions. The reference in this case was the IRs of an FDN of size $N=64$, ensuring a sufficiently smooth response. The frequency-dependent attenuation filter was designed following the steps outlined in Section 4.1.
The test evaluated the coloration of the standard differentiable FDN (DiffFDN) proposed in this paper, the DiffFDN with scattering matrix (SCAT), and that with optimizable Householder matrix (HH). Each configuration was tested on a separate page where the number and lengths of the delays were constant, and only the feedback matrix and the input and output gains were altered. The FDN implementation of the Schroeder series allpass reverberator (SH) was the anchor, whereas the random orthogonal feedback matrix (RO) acted as a baseline. For the RO condition, the initial values of optimization of the DiffFDN were used. The direct gain $d$ was set to zero in all cases. In total, there were 18 listening test pages with six stimuli each. The loudness of each IR was normalized to ensure consistency across conditions.
The experiment was conducted in a sound-insulated booth at the Aalto Acoustics Lab, with participants wearing Sennheiser HD 650 headphones. The final items were presented to 14 listeners. The average age of the participants analyzed was 29.5 years, with a standard deviation of 4.72, and none of them reported any hearing impairments. All participants, except one, were either students or employees of the Aalto University Acoustics Lab and had previous experience with the MUSHRA test.
The box plots presented in Fig. 6 illustrate the outcomes of the listening test. In each box, the median is represented by the central mark, while the lower and upper edges indicate the 25th and 75th percentiles, respectively. The whiskers extend to encompass the most extreme data points not identified as outliers, with any outliers plotted separately. The shaded regions surrounding the medians facilitate the comparison of sample medians across various box charts. Non-overlapping shaded regions signify differing medians between the compared box charts at the 5% significance level, assuming a normal distribution.
Despite excluding the reference and anchor conditions, the data, as revealed by the Shapiro-Wilk test [31], deviated from a normal distribution. Additionally, the Wilcoxon signed-rank test [32] was employed to assess the score distributions for each pair of conditions within each page. To address multiple comparisons (15 hypotheses per page), we applied the Bonferroni method to adjust the alpha level.
The $p$-values for most paired conditions indicate significant differences among all pairs of results. However, there are a few exceptions observed in the distribution of DiffFDN and HH. In these cases, the $p$-values exceed the alpha level across all conditions with decaying modes, suggesting a convergence in the scores assigned to the two conditions. Moreover, DiffFDN and HH reject the null hypothesis when paired with SCAT in the lossy cases for $N=8$, suggesting that for larger FDN sizes, the performance improvement provided by SCAT diminishes.
The IRs of the optimized FDNs always show a median greater than the initial condition (RO), proving the effectiveness of the optimization. In the late reverb analysis, Fig. 6 (a), the DiffFDN has median ratings of 74, 82.5, and 70, respectively, for increasing FDN size. Only for $N=$ 8, the FDN with scattering matrix received ratings higher than DiffFDN with a median of 76.
Introducing early reflections provides an advantage for the FDN with a scattering matrix (SCAT) compared to other conditions. In general, spectral flatness is associated with temporal smoothness. The scattering feedback matrix increases echo density during early reflections, potentially suggesting a less colored response, as shown in Fig. 6 (b) and (c). This effect becomes more noticeable for smaller FDNs, and for $N=8$, the medians of the results begin to converge among optimized FDNs.
The confidence intervals for the results of all tests are relatively large. Although the objective results in Table 4 indicate a limited reduction in modal excitation standard deviation, the optimized Householder matrix was rated relatively close to the DiffFDN in the listening test. The distribution of modal excitation is one quantity that can aid in deducing the coloration of the reverberator. However, this is a much more complicated concept, as evidenced by the results presented in the listening test.
7 Conclusion
The presented work introduces an optimization method for designing artificial reverberation using a differentiable feedback delay network (DiffFDN). The method focuses on achieving spectral flatness and temporal density by optimizing the feedback matrix, as well as the input and output gains. Emphasis has been placed on improving the computational efficiency of the method without compromising its performance. By incorporating attenuation filters and an optimizable scattering feedback matrix, this method can be further enhanced, presenting an efficient option for synthesizing natural room impulse responses.
Favorable results were obtained in this study with a DiffFDN employing as few as four delay lines. In objective evaluations, we demonstrated that the proposed method reduces the width of the modal excitation distribution, thereby decreasing the prominence of the loudest modes. Listening test results further confirmed the success of this approach in attenuating artifacts commonly associated with undesired coloration, making the DiffFDN close to an ideally colorless, smooth reverberator.
Declarations
Availability of data and materials
The PyTorch implementation of the proposed method can be found on the online repository  [^1] https://github.com/gdalsanto/diff-fdn-colorless/. Listening test stimuli examples can be found on the dedicated website  [^2] http://research.spa.aalto.fi/publications/papers/eurasip-colorless-fdn/.
Competing interests
The authors declare that they have no competing interests.
Funding
The Aalto University School of Electrical Engineering funded the work of the first author.
Authors’ contribution
GDS performed the methodology, analysis, validation, and wrote the original draft. KP performed the conceptualization, writing, and project supervision. SJS performed conceptualization, methodology, analysis, writing, and project supervision. VV performed project supervision, review, and editing.
Acknowledgments
Not applicable

### v2

@@
-Feedback Delay Network Optimization
+Efficient Optimization of Feedback Delay Networks for Smooth Reverberation
 Abstract
@@
 The proposed method achieves computational savings compared to the baseline while preserving its performance.
-The effectiveness of this work is demonstrated through two application scenarios where natural-sounding synthetic impulse responses are obtained via the introduction of attenuation filters and an optimizable scattering feedback matrix.
+The effectiveness of this work is demonstrated through two application scenarios where smooth-sounding synthetic impulse responses are obtained via the introduction of attenuation filters and an optimizable scattering feedback matrix.
 keywords:
@@
 However, these efforts proved insufficient, necessitating additional considerations regarding the selection of parameters to effectively reduce or eliminate this coloration effect [4, 5, 2, 6].
-Recently, the perceived coloration was attributed to the modal properties of synthesized reverbs, with a wide distribution of modal excitations correlating positively with increased coloration [7].
-Therefore, modal decomposition of artificial reverberation architectures, such as FDNs [8], can aid in identifying and mitigating phenomena that are linked to perceptual artifacts.
+Recently, the perceived coloration was attributed to the modal properties of synthesized reverbs, with a wide distribution of modal excitations linked to increased coloration [7].
+Therefore, modal decomposition of artificial reverberation architectures, such as FDNs [8], can aid in identifying phenomena that are related to perceptual artifacts.
 In FDNs, the values of the modal excitations depend on all system parameters [8].
 In this paper, the objective is to optimize the gain parameters and feedback matrix of an FDN.
-The goal is to minimize the perceptual coloration observed in its impulse response (IR).
-We adopt a stochastic gradient descent approach, focusing specifically on the magnitude of the frequency response.
-This paper extends our previous work [9] in which part of the loss function was based on the IR.
-The current work also demonstrates two practical applications of the proposed approach in synthesizing natural-sounding IRs by including attenuation filters in the FDN structure and optimizing the scattering feedback matrix.
-As we now concentrate solely on the frequency domain, the computational requirements are streamlined.
-We attain control over temporal density by utilizing its correlation with the density of the feedback matrix entries.
+Similarly to our previous work [9], the goal is to minimize the perceptual coloration observed in the impulse response (IR) by adjusting the parameters using stochastic gradient descent.
+In [9], we targeted the flatness of the magnitude response and temporal density by incorporating two loss functions in the frequency and time domains, respectively.
+For the time domain, the inverse discrete Fourier transform (DFT) operation was needed, which presented a bottleneck that we aim to address in this paper.
+The current work also demonstrates practical applications of the proposed approach in synthesizing natural-sounding late reverberation by optimizing the scattering feedback matrix and including in the FDN structure attenuation filters that benefit from the optimization.
+As we now concentrate solely on the frequency domain, the computational requirements are streamlined by analyzing the frequency response sparsely using batch processing over subsets of frequency samples.
+We attain control over temporal density by utilizing its correlation with the density of the feedback matrix entries instead of computing the system’s IR.
 Evaluation is conducted by analyzing the modal excitation distribution.
 A perceptual evaluation against the FDN design before optimization shows that the proposed method successfully decreased perceived coloration.
+This work distinguishes itself from a recent and related study presented in [10] by addressing more fundamental issues underlying the limitations in reverb synthesis using FDNs.
+At the same time, it suggests how this approach can be extended to IR synthesis tasks by introducing a filter design method of our choice.
+Nonetheless, the main contribution of the presented paper does not depend on the filter used, thereby opening a wider range of applications.
 The paper is structured as follows.
@@
 2.1 Feedback delay network
-Figure 1: FDN with a single input, a single output, and an $N$-by-$N$ feedback matrix A. The thick lines indicate $N$-channel signal paths. The parameters under optimization are highlighted in blue.
-An FDN is a recursive system consisting of delay lines, a set of gains, and a scalar feedback matrix through which the delay outputs are coupled to the delay inputs.
+Figure 1: FDN with a single input, a single output, and an $N$-by-$N$ feedback matrix A.
+The thick lines indicate $N$-channel signal paths.
+The circled multiplication symbols indicate a vector multiplication.
+The parameters under optimization are highlighted in blue.
+A FDN is a recursive system consisting of delay lines, a set of gains, and a scalar feedback matrix through which the delay outputs are coupled to the delay inputs.
 Fig.
@@
 2.2 Modal decomposition
-The IR of the FDN can be expressed as the sum of complex one-pole modes, also known as resonators [8]:
-$h(n)=\sum_{i=1}^{\mathcal{{M}}}h_{i}(n)\;.$ (4)
-Each mode $h_{i}(n)$ is defined by the pole $\lambda_{i}$ and the residue $\rho_{i}$:
-$h_{i}(n)=|\rho_{i}||\lambda_{i}|^{n}e^{\jmath(n\angle\lambda_{i}+\angle\rho_{i})}\,,$ (5)
+The IR of the FDN can be expressed as the sum of complex one-pole modes, also known as resonators, each defined by its pole $\lambda_{i}$ and its residue $\rho_{i}$ [8]:
+$h(n)=\sum_{i=1}^{\mathcal{{M}}}|\rho_{i}||\lambda_{i}|^{n}e^{\jmath(n\angle\lambda_{i}+\angle\rho_{i})}\;.$
 where ||$\cdot$|| ⋅ | denotes the magnitude, $\angle$ represents the argument of a complex number in radians, $\jmath=\sqrt{-1}$, and $n$ indicates the discrete time index.
@@
 The magnitude and phase of the $i^{\text{th}}$ pole, $|\lambda_{i}|$ and $\angle\lambda_{i}$, respectively, dictate its decay rate and frequency.
-A common approach to calculate the modal decomposition (6) from the rational polynomial form of the transfer function (3) involves determining the system poles $\lambda_{i}$.
-These poles are obtained as the roots of the GCP (2), derived from the eigenvalues of the linearized feedback matrix [8].
+A common approach to calculate the modal decomposition (4) from the rational polynomial form of the transfer function (3) involves determining the system poles $\lambda_{i}$ from the eigenvalues of the linearized feedback matrix [8].
+An important aspect for the remainder of this paper is the distribution of the modal excitation values $|\rho_{i}|$.
+While the decay of the modes $|\lambda_{i}|$ is typically governed entirely by the target reverberation time $T_{60}$, the modal excitation remains largely unconstrained by design.
+The connection between the distribution of modal excitation with perceived coloration [7] is a key finding underlying the proposed optimization method
 2.3 Homogeneous decay in FDN
 Designing an artificial reverberator with FDN often starts by creating a lossless prototype, having an energy-preserving feedback loop [11, 12].
-The FDN is said to be lossless if the roots of the GCP (2) have magnitude equal to one, i.e., $|\lambda_{i}|=1$ for all $i$ [12].
+The FDN is said to be lossless if the roots of the GCP (2) have magnitude equal to one, i.e., $|\lambda_{i}|=1$ for all $i\,$s [13].
 The advantage of initially designing a lossless FDN lies in the straightforward implementation of frequency-dependent decay that equally influences all system poles.
@@
 In this study, we mainly focus on the specific case of frequency-independent homogeneous decay.
-This refers to the case where all modes experience the same decay rate, i.e., $|\lambda_{i}|=\gamma$ for all $i$’s.
-In Subsection 4.1, we will demonstrate how an FDN optimized at such $\gamma$ can be employed as a lossless prototype for synthesizing reverb with frequency-dependent decay.
-Homogeneous decay is achieved with a feedback matrix $\bm{A}$ being the product of a unilossless matrix $\bm{U}$ and a diagonal matrix $\bm{\Gamma}$, whose entries are delay-proportional absorption coefficients, $\bm{\Gamma}=\text{diag}(\gamma^{\bm{m}})$.
+This refers to the case where all modes experience the same decay rate, i.e., $|\lambda_{i}|=\gamma$ for all $i$ s.
+In Subsection 4.1, we will demonstrate how an FDN optimized as such can be employed as a lossless prototype for synthesizing reverb with frequency-dependent decay.
+Homogeneous decay is achieved with a feedback matrix $\bm{A}$ being the product of a unilossless matrix $\bm{U}$ and a diagonal matrix $\bm{\Gamma}$, whose entries are delay-proportional absorption coefficients, $\bm{\Gamma}=\text{diag}(\gamma^{\bm{m}})$, $\gamma^{\bm{m}}$ represents the $\gamma$ value raised to the power of each corresponding delay line in $\bm{m}$.
 The feedback matrix can be expressed as
@@
 $\gamma_{\mathrm{dB}}=\frac{-60}{f_{\textrm{s}}T_{60}}\;,$ (6)
-where $f_{\textrm{s}}$ is the sampling rate in Hz and $T_{60}$ is the reverberation time in s defined as the time required for the sound level to decay by $60$ dB from the initial steady-state value.
+where $f_{\textrm{s}}$ is the sampling rate in Hz.
 2.4 Coloration in FDN
-Figure 2: Magnitude response of an FDN with size $N=8$, showcasing the difference between long and short delay lines when the former is 10 times longer than the latter.
-Natural late reverberation can be conceptualized as white noise with an exponential decay, providing perceptually ideal smooth diffuse reverberation [14].
+Figure 2: Magnitude response of a homogeneous FDN with size $N=8$, diagonal matrix, and unit input and output gains, showcasing the difference between long and short delay lines when the former is 10 times longer than the latter.
+Natural late reverberation can be modeled as white noise with an exponential decay, providing perceptually ideal smooth diffuse reverberation [15].
 In artificial reverberation, achieving a white noise generator involves ensuring the allpass property is maintained.
@@
 Listening test results suggest that a narrow distribution of the modal excitation values $|\rho_{i}|$ tends to result in a perceptually white spectrum [7].
-For large values of $|\rho_{i}|$, coloration starts to become noticeable [7].
-We also observed this outcome in our earlier study, where we conducted a listening test focusing on the coloration of FDN-based reverb [9].
-In this test, we compared the reverb before and after applying an optimization method to narrow the modal excitation distribution.
-The literature also indicates that more than 6000 modes are needed for an IR to be perceived as rather colorless [15].
+However, when a subset of modes exhibit high $|\rho_{i}|$ values relative to the mean of the system’s modal excitation distribution, coloration starts to become noticeable [7].
+We also observed this outcome in our earlier study, where we conducted a listening test focusing on the coloration of a FDN-based reverb [9].
+In this test, we compared the reverb before and after applying an optimization method, which resulted in a narrower modal excitation distribution.
+The literature indicates that more than 6000 modes are needed for an IR to be perceived as rather colorless [16].
 Additionally, it was shown that for large values of $\mathcal{{M}}$, the modes of the FDNs are uniformly distributed [8], preventing additional coloration that usually results from clusters of modes.
@@
 Specifically, when the feedback matrix $\bm{A}$ is diagonal, the FDN adopts the configuration of a parallel comb-filter structure.
-In the case of a homogeneous FDN with a diagonal feedback matrix, the transfer function in (1) is analogous to a superposition of comb filters, with each filter having the transfer function
-$H_{i}(z)=\frac{1}{1+\gamma^{m_{i}}z^{-m_{i}}}\,.$ (9)
+In the case of a homogeneous FDN with a diagonal feedback matrix, the transfer function in (1) is analogous to a superposition of comb filters with the delay line in the feedforward path and the gain $\gamma^{m}_{i}$ in the feedback.
+Each filter has a transfer function
+$H_{\textrm{comb}_{i}}(z)=\frac{z^{-m_{i}}}{1-\gamma^{m_{i}}z^{-m_{i}}}\,.$ (7)
 The contribution of each filter to the total energy of the response can be calculated as
-$\lVert H_{i}(z)\rVert_{2}^{2}$ $=\int_{0}^{2\pi}\lvert H_{i}(e^{\jmath\omega})\rvert^{2}\textrm{d}\omega$ (10)
+$\lVert H_{\textrm{comb}_{i}}(z)\rVert_{2}^{2}$ $=\int_{0}^{2\pi}\lvert H_{\textrm{comb}_{i}}(e^{\jmath\omega})\rvert^{2}\textrm{d}\omega$ (8)
 $=\frac{1}{1-\gamma^{2m_{i}}}\,,$ (9)
@@
 In contrast, longer delays contribute less energy and tend to be masked by the more dominant comb filters [17].
+The same conclusion can be derived through pole-zero analysis of (7).
+By definition, FDNs are networks of comb filters whose feedback paths are interconnected through a feedback matrix.
+Therefore, the reasoning above applies to FDNs with dense feedback matrices without loss of generality.
 Figure 2 shows a section of the magnitude spectrum of two FDNs with identical parameters but different delay-line lengths.
+Since the resonant frequencies are uniformly distributed across the spectrum, [8] the choice of the observed frequency range is arbitrary and serves only for visualization purposes.
+The feedback matrix is diagonal with values alternating between +1 and -1.
+The input and output gains are unit vectors.
 The FDN with shorter delay lengths, approximately one-tenth of the longer delays, exhibits higher magnitude values and more distinct, louder resonances.
 For long delays, the resonances become less noticeable, and the loudest frequencies exhibit a magnitude approximately 20 dB lower compared to the scenario with shorter delays.
-This example shows that to achieve a colorless FDN, it is necessary to avoid strongly recirculating short delays and encourage exciting long delays instead.
+This example suggests that to achieve a colorless FDN, it is necessary to avoid strongly recirculating short delays and instead focus on exciting long delays.
 2.5 Problem statement
@@
 1, the FDN parameters being optimized are indicated in blue.
-Previous studies indicate that the choice of frequency-dependent attenuation has little impact on coloration [7].
-Therefore, we conduct optimization on a long-ringing, frequency-independent prototype FDN.
-For ideal late reverberator synthesis with this setup, the spectrum’s magnitude must remain constant across frequencies and time.
+Previous studies indicate that the choice of frequency-dependent attenuation has little impact on coloration [7] Therefore, we conduct optimization on a long-ringing, frequency-independent prototype FDN.
+For ideal late reverberator synthesis with this setup, the spectrum’s magnitude must remain constant across frequency and time.
 However, the recursive structure of FDN introduces undesired colorations, which constitutes a primary challenge in our work, as emphasized in Subsection 2.4.
@@
 In this section, we present a method to reduce coloration in the IR of an FDN with frequency-independent $T_{60}$.
-We employ stochastic gradient descent to optimize the parameters of the FDN, which has been made differentiable by utilizing the frequency sampling method.
+We employ stochastic gradient descent to optimize the gain parameters of the FDN, which has been made differentiable by utilizing the frequency sampling method.
 3.1 Differentiable FDN
@@
 An FIR filter is straightforward to implement, as it does not require a recursive structure.
-Moreover, an FIR filter can be easily parallelized and optimized for hardware acceleration, such as a graphics processing unit (GPU).
+Moreover, an FIR filter can be easily parallelized and optimized for hardware acceleration, such as a graphics processing unit.
 To obtain the FIR approximation of $H(z)$, the delay matrix $\bm{D_{m}}(z)$ is evaluated at discrete frequency points
 […]
-where $M$ indicates the total number of frequency bins equally distributed on the unit circle.
+where $M$ indicates the total number of frequency bins evenly distributed on the unit circle.
 The block diagram of the architecture is depicted in Fig.
@@
 During training, the FDN is set to have a homogeneous decay by forcing $\bm{A}$ to satisfy (5) for a given $\gamma$.
-Figure 3: Architecture of the proposed optimization workflow. A thick line indicates a $N$-channel signal path. Backpropagation is highlighted in blue.
+Figure 3: Architecture of the proposed optimization workflow. A thick line indicates an $N$-channel signal path. Backpropagation is highlighted with a dashed blue line.
 At each training step the estimated channel-wise transfer function $\bm{H}(z)$ is computed at a batch of frequency bins,
@@
 Every batch is composed of $\mu<M$ different frequency points that are randomly sampled from $\bm{z}_{M}$.
-The value of $M$ is such that to ensure oversampling, while we found empirically that $\mu$ can be as small as $0.05f_{\textrm{s}}$.
+The value of $M$ is such that to ensure oversampling, we found empirically that $\mu$ can be as small as $0.05f_{\textrm{s}}$.
 At each epoch, $\bm{H}(z)$ is estimated at all frequency points within the vector $\bm{z}_{M}$.
@@
 To guarantee frequency independence of the energy decay, we adopted the parameterization in (5), where the feedback matrix is computed from an orthogonal matrix $\bm{U}$.
-The class of orthogonal matrices meets the unitary condition for losslessness and can be obtained through a parameterization of skew-symmetric matrices.
+The class of orthogonal matrices meets the unitary condition for losslessness and can be obtained through the parameterization of skew-symmetric matrices described in [18].
 At initialization, a random matrix $\bm{W}$ is constructed. At each optimization step, $\bm{W}$ is mapped to a skew-symmetric matrix, and the matrix exponential is computed,
@@
 3.3 Gain-per-sample
-Figure 4: Magnitude response of a homogeneous FDN of order $\mathcal{M}=8944$ with random orthogonal feedback matrix and unitary input/output gains at different values of $T_{60}$ in seconds.
+Figure 4: Magnitude response of a homogeneous FDN of size $N=8$ and order $\mathcal{M}=8944$ with random orthogonal feedback matrix and unitary input/output gains at different values of $T_{60}$ in seconds.
 In the lossless case, $T_{60}=\infty$, the resonances tend to infinity.
@@
 For $T_{60}=0.14\,$s ($\gamma=0.999$), the resonance half-widths become excessively large, making it difficult to identify individual resonances.
+While (6) provides a lower bound for $\gamma$, defining an upper bound is more complex.
+Therefore, we recommend treating it as a hyperparameter to be tuned for optimal performance.
 We ran optimization at different values of $\gamma$, and we observed good convergence when $T_{60}\leq 10$ s.
@@
 In stochastic gradient descent, as in many other optimization algorithms, the choice of initial parameter values can affect the algorithm’s ability to find the global or a good local minimum of the loss function.
-To start from a point that facilitates an efficient exploration of the loss landscape, we initialize the values of $\bm{b}$ and $\bm{c}$ by drawing them from the normal distribution $\mathcal{N}(0,N^{-1})$, and the entries of $\bm{W}$ from the uniform distribution $\mathcal{U}(-1/\sqrt{N},1/\sqrt{N})$.
+We initialize the values of $\bm{b}$ and $\bm{c}$ by drawing them from the normal distribution $\mathcal{N}(0,N^{-1})$, and the entries of $\bm{W}$ from the uniform distribution $\mathcal{U}(-1/\sqrt{N},1/\sqrt{N})$.
+This initialization strategy has empirically yielded good results in our experiments.
 The design of the delays is a rather non-trivial task that requires further constraints.
@@
 To avoid degenerative patterns and ensure a smooth-sounding reverb, we choose delays that are logarithmically distributed co-prime numbers leading to $\mathcal{{M}}\geq 6000$.
+The design of delays depends on the specific application, and the gain parameters must be optimized whenever a new set of delay lines is used since they contribute to the system’s modal excitation distribution.
 3.5 Loss function
-Figure 5: Evolution of losses (left) and the relative density increase of the IR (right).
+Figure 5: Evolution of losses (left) during training of a size 8 FDN with an orthogonal feedback matrix trained with $\alpha=1$ (blue) and $\alpha=0$ (green).
+On the right, the echo density profile [21] at initialization and after optimization for both cases.
 The network is trained on two losses, $\mathcal{L}_{\textrm{spectral}}$ and $\mathcal{L}_{\textrm{sparsity}}$.
@@
 The total loss function is
-$\mathcal{L}=\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))+\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ (18)
-and the individual loss terms are
-$\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))$ $=\frac{1}{\mu}\sum_{z\in\bm{z}_{\mu}}\sum_{i=1}^{N}\left({\lvert H_{i}(z)\rvert-1}\right)^{2}$
+$\mathcal{L}=\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))+\alpha\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ (16)
+where $\alpha$ is a weighting term that controls the influence of $\mathcal{L}_{\textrm{sparsity}}$ on the overall loss. In this work, we set $\alpha$ to 1. The individual loss terms are
+$\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))$ $=\frac{1}{\mu}\sum_{z\in\bm{z}_{\mu}}\frac{1}{N}\sum_{i=1}^{N}\left({\lvert H_{i}(z)\rvert-1}\right)^{2}$ (17)
 $+\left({\lvert H(z)\rvert-1}\right)^{2}$
-$\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ $=\frac{N\sqrt{N}-\sum_{i,j}\lvert U_{ij}\rvert}{N(\sqrt{N}-1)}$
+$\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ $=\frac{N\sqrt{N}-\sum_{i,j}\lvert U_{ij}\rvert}{N(\sqrt{N}-1)},$ (18)
 where $H_{i}(z)$ is the output of the network’s $i^{\text{th}}$ channel computed from the output of the $i^{\text{th}}$ delay line and scaled by $c_{i}$.
+Including the channel-wise loss term in (17) is optional, as our tests showed comparable results on the objective metrics with and without it.
 The sparsity loss, which is the last term above, is calculated from the sum of the absolute values of the matrix entries, i.e.
 $\sum_{i,j}\lvert U_{ij}\rvert$.
-For orthogonal matrices, its value is bounded by $-N\sqrt{N}$ and $N$.
-To make the dynamic range of the $\mathcal{L}_{\textrm{sparsity}}$ comparable with that of $\mathcal{L}_{\textrm{spectral}}$ we shifted $\sum_{i,j}\lvert U_{ij}\rvert$ by $N$ and normalized it by $N(\sqrt{N}-1)$.
-The term is then shifted by $1$ to span positive values only.
+For orthogonal matrices, its value is bounded between $N$ and $N\sqrt{N}$.
+To make the dynamic range of the $\mathcal{L}_{\textrm{sparsity}}$ comparable with that of $\mathcal{L}_{\textrm{spectral}}$ we shifted $\sum_{i,j}\lvert U_{ij}\rvert$ by $N\sqrt{N}$ and normalized it by $N(\sqrt{N}-1)$.
+This sparsity loss is motivated by the fact that echo density increases when the delay lines are highly interconnected.
+This principle underlies the FDN’s structure, distinguishing it from that of parallel comb-filters [14, 3].
+Moreover, in our previous work [9] we found that the absence of a term controlling the temporal density of the system’s IR may lead the matrix $\bm{U}$ to converge towards either a diagonal matrix or its permutation.
 In our previous work [9], the sparseness was controlled by a temporal loss consisting in the ratio of the $\ell_{1}$ norm to the $\ell_{2}$ norm of the estimated IR.
-This required the inverse Fourier transform, slowing down the optimization steps.
-To visualize the impact the loss function (18) has on the IR density, Fig.
-5 illustrates how the $\ell_{1}$ norm of the estimated IR signal $h$, i.e., $\sum_{k}\lvert h(k)\rvert$, evolves in relation to the losses at each epoch of one run of training.
-The right pane shows the increase in density relative to initialization.
-The left pane shows the evolution of the loss and its terms at each epoch.
-The terms decay at similar rates, and after 10 epochs, they start to converge.
-Compared to the losses, the increase in density starts to converge quicker, reaching a $16\%$ increase after just 5 epochs.
-Without sparsity control, the magnitude response may converge towards that of a comb filter, leading to a corresponding decrease in IR density, as demonstrated in our previous work [9].
+This required the inverse DFT, slowing down the optimization steps.
+To clarify the impact of the loss function (16) on echo density, Fig.
+5 illustrates how echo density profile, as defined in [21], is influenced by the proposed training, particularly by the inclusion of $\mathcal{L}_{\textrm{sparsity}}$.
+The left panel in Fig.
+5 displays the evolution of the loss and its components over each epoch.
+The plot shows the losses when an instance of an FDN with eight delay lines is trained using (16) with $\alpha=1$, as well as using only $\mathcal{L}_{\textrm{spectral}}$, i.e.
+$\alpha=0$ in (16), indicated in the plot with an asterisk.
+In Fig.
+5 (left), the terms decay at similar rates, and after six epochs, they begin to converge.
+The right panel presents the echo density of the FDN at initialization, and after optimization with $\alpha=1$ and $\alpha=0$.
+Both optimizations demonstrate an improvement in echo density, as the curve reaches 1 faster.
+However, when $\mathcal{L}_{\textrm{sparsity}}$ is included (i.e., $\alpha=1$), the echo buildup occurs more rapidly from the early reflections.
+Moreover, without sparsity control, the magnitude response may converge towards that of a comb filter, leading to a corresponding decrease in IR density, as demonstrated in our previous work [9].
+To implement the differentiable model and the training iterations, we used the PyTorch library [22].
+PyTorch handles parameter adjustments through backpropagation, which involves computing the gradient of the loss function with respect to the given parameters.
+PyTorch’s built-in differentiation engine automatically computes the gradients for any computational graph.
+Additionally, PyTorch supports complex-valued tensors and backpropagation for real-valued functions of complex tensors, using Wirtinger calculus, which is particularly beneficial for our work in the frequency domain [23].
 4 Applications
@@
 Using a differentiable FDN optimized for colorlessness as a lossless prototype can address coloration in the prototype that cannot be effectively mitigated by the attenuation filter, facilitating a reduction in size as well.
-The attenuation filter is designed to approximate a target frequency-dependent reverberation time $T_{60}(\omega)$ by achieving a target gain-per-sample $\gamma(\omega)$
+The attenuation filter is designed to approximate a target frequency-dependent reverberation time $T_{60}(\omega)$ by achieving a target gain-per-sample $\gamma(\omega)$:
 $\gamma_{\textrm{dB}}(\omega)=-\frac{60}{f_{\textrm{s}}T_{60}(\omega)}\,.$ (19)
 Attenuation filters are commonly introduced after the delay lines [24].
-To simplify the transfer function, at the price of a negligibly lower $T_{60}(\omega)$ accuracy, we place the filters in the feedback loop, resulting in the following transfer function
+To simplify the transfer function, at the price of a negligibly lower $T_{60}(\omega)$ accuracy, we place the filters in the feedback loop, resulting in the following transfer function:
 $H(z)=\bm{c}^{\top}\big{[}\bm{D_{m}}(z)^{-1}-\bm{U}\bm{\Gamma}(z)\big{]}^{-1}\bm{b}+d,$ (20)
@@
 When using the Householder matrix $\bm{U}=\bm{I}-2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}$, where $\bm{v}$ is a unit vector, gives a simple, yet restricted, parametrization of orthogonal matrices.
-Using the Sherman-Morrison formula, i.e.,
-(${\lparen\bm{W}+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}\rparen}^{\text{-}1}={\bm{W}}^{\text{-}1}-\frac{2{\bm{W}}^{\text{-}1}\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{W}}^{\text{-}1}}{1+2\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{W}}^{\text{-}1}\bm{v}}$ (25)
+Using the Sherman-Morrison formula for an invertible square matrix $\bm{V}$, i.e.,
+(${\lparen\bm{V}+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}\rparen}^{\text{-}1}={\bm{V}}^{\text{-}1}-\frac{2{\bm{V}}^{\text{-}1}\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{V}}^{\text{-}1}}{1+2\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{V}}^{\text{-}1}\bm{v}}$ (25)
 can help to avoid the matrix inverse. Thus, the FDN transfer function is then written as
-$H(z)=\bm{c}^{\top}{\left({\bm{W}(z)+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}}\right)}^{\text{-}1}\bm{b}\,,$ (26)
-where $\bm{W}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{I}$.
+$H(z)=\bm{c}^{\top}{\left({\bm{V}(z)+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}}\right)}^{\text{-}1}\bm{b}\,,$ (26)
+where $\bm{V}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{I}$.
 With this parameterization, the computational complexity required to calculate the FDN transfer function at one frequency point is reduced from $\mathcal{O}(N^{3})$ to $\mathcal{O}(N)$.
 To ensure homogeneous absorption, $\bm{\Gamma}$ is included in (26) as
-$H(z)=\bm{c}^{\top}{\left({\bm{W_{\gamma}}(z)+2\bm{v}\bm{v_{\gamma}^{\mathstrut\scriptscriptstyle\top}}}\right)}^{\text{-}1}\bm{b}\,,$ (27)
-where $\bm{W_{\gamma}}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{\Gamma}$ and $\bm{v_{\gamma}}^{\mathstrut\scriptscriptstyle\top}=\bm{v}^{\mathstrut\scriptscriptstyle\top}\Gamma$.
+$H(z)=\bm{c}^{\top}{\left({\bm{V_{\gamma}}(z)+2\bm{v}\bm{v_{\gamma}^{\mathstrut\scriptscriptstyle\top}}}\right)}^{\text{-}1}\bm{b}\,,$ (27)
+where $\bm{V_{\gamma}}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{\Gamma}$ and $\bm{v_{\gamma}}^{\mathstrut\scriptscriptstyle\top}=\bm{v}^{\mathstrut\scriptscriptstyle\top}\Gamma$.
 With this parametrization, the optimizable parameters become $\bm{v},\,\bm{b},\,\text{and}\,\bm{c}$.
@@
 The delay lengths are prime numbers distributed logarithmically.
-In all configurations, the total number of modes is $90008700<\mathcal{M}<$ 9000.
-Table 1: Delay-line lengths for each size $N$ of the analyzed FDNs. Values are logarithmically distributed prime numbers.
+In all configurations, the total number of modes is $90008700<\mathcal{M}<$ 9000, which ensures us to rule out artifacts related to low modal density [16].
+Table 1: Delay-line lengths (in samples) for each size $N$ of the analyzed FDNs. The lengths are logarithmically distributed prime numbers.
 N Lengths
@@
 […]
-During optimization, we used a sampling rate of $f_{\textrm{s}}=48$ kHz and $M=480000$ frequency points equally distributed over $[0,\pi]$.
+During optimization, we used a sampling rate of $f_{\textrm{s}}=48$ kHz and $M=480000$ frequency points evenly distributed over $[0,\pi]$.
 The batch size was $\mu=2000$, requiring 240 steps to complete an epoch.
@@
 This choice implies $T_{\textrm{60}}=1.439\,$s and satisfies (15).
-The configurations under analysis include the differentiable FDN optimized according to our prior work (DAFx23) [9], the one optimized with the proposed method (DiffFDN), and two variations of the latter: one with an optimizable Householder matrix (HH) and another with an optimizable scattering matrix (SCAT).
+The configurations under analysis include the differentiable FDN optimized according to our prior work (DAFx23) [9], which employs an orthogonal feedback matrix, as well as the one optimized with the proposed method (DiffFDN-O) that also features an orthogonal feedback matrix, and two variations of the latter: one with an optimizable Householder matrix (DiffFDN-HH) and another with an optimizable scattering matrix (DiffFDN-SCAT).
 5.2 Computational complexity
@@
 During the forward pass, the computational complexity needed to calculate the FDN transfer function at a single frequency point in (11) is $\mathcal{O}(N^{3})$.
-For each epoch, the proposed method (DiffFDN) computes the transfer function at $M$ points.
-In contrast, the baseline approach (DAFx23) requires a similar number of frequency points at each training step, resulting in a linear increase in the number of operations with the dataset size, e.g., 256 in [9].
+For each epoch, the proposed method (DiffFDN-O) computes the transfer function at $M$ points, where $M$ corresponds to the dataset size.
+In contrast, the baseline approach (DAFx23) requires calculating a similar number of frequency points at each training step, leading to a linear increase in the number of operations as the dataset size grows, which in [9] it is represented by different values of $M$ (e.g., 256).
 This is because the number of epochs required for convergence in the baseline is close to that of the presented method.
-Additionally, in the forward pass of the baseline, the FFT is necessary due to the temporal loss, introducing additional operations.
+Additionally, in the forward pass of the baseline, the DFT is necessary due to the temporal loss, introducing additional operations.
 By employing the Householder matrix, we can significantly decrease the computational complexity of (11) to $\mathcal{O}(N)$.
-Introducing the FFT would raise this complexity to $\mathcal{O}(N\text{log}_{2}(N))$.
+Introducing the DFT would raise this complexity to $\mathcal{O}(N\text{log}_{2}(N))$.
 The FDNs also have different computational complexity during operation.
@@
 RO 208 324 448
-DiffFDN 208 324 448
-DiffFDN (HH) 200 300 400
-DiffFDN (SCAT) 288 480 704
+DiffFDN-O 208 324 448
+DiffFDN-HH 200 300 400
+DiffFDN-SCAT 288 480 704
 5.3 Modal excitation distribution
@@
 Based on this test alone, it is challenging to predict the overall trend across different FDN sizes.
-Table 3: Standard deviation of the modal excitation before (Init) and after optimization (Optim) at different FDN sizes.
+Table 3: Standard deviation of the modal excitation of DiffFDN-O before (Init) and after optimization (Optim) at different FDN sizes.
 N Init Optim
@@
 […]
-To ensure that the changes to the optimization framework did not reduce the performance of the method introduced in [9], we conducted identical tests using our previous optimization framework.
+To assess the impact of the changes to our previous optimization framework [9]—particularly the addition of the sparsity loss function and sparse frequency sampling—we ran the same tests with the original optimization framework for comparison.
 Each set of parameters used at initialization is shared between the two methods.
@@
 Even though it is likely impossible to perceive a difference of 0.03 dB in standard deviation, the more significant improvement lies in the reduction of operation count.
-Table 4 also shows the results of a similar test run on the FDN with optimizable scattering matrix (SCAT) and Householder matrix (HH), both of size $N=4$.
-Due to a different parameterization of the feedback matrix, the distributions of the initial parameters are different from the standard configuration.
+Table 4 also shows the results of a similar test run on DiffFDN-SCAT and DiffFDN-HH, both of size $N=4$.
+Due to a different parameterization of the feedback matrix, the distributions of the initial parameters are different from both DAFx23 and DiffFDN-O.
 After optimization, the standard deviation reduces in both cases.
-Among all tested configurations, the HH case exhibited the smallest improvement, likely due to its more restrictive parameterization.
-The SCAT configuration also showed a lower degree of improvement compared to the standard configuration.
+Among all tested configurations, the DiffFDN-HH case exhibited the smallest improvement, likely due to its more restrictive parameterization.
+The DiffFDN-SCAT case also showed a lower degree of improvement compared to DiffFDN-O.
 This could be attributed to the higher complexity of the problem, suggesting potential enhancement through fine-tuning the optimization framework.
-Table 4: Standard deviation of the modal excitation before (Init) and after optimization (Optim) of different frameworks, each employing an FDN of size N=4 and identical set of dealys.
+The histograms in Fig.
+6 illustrate the distribution of modal excitation at the start (Init) and end of the optimization process (Optim) for a single instance of DiffFDN-O, across all analyzed sizes.
+The modal excitation values have been centered around 0 dB.
+The optimization reduces the loudest modes, leading to a shift toward a narrower overall excitation distribution.
+Figure 7 shows the effect on their magnitude spectrum.
+Similarly to Fig.
+2, only a narrow section of the spectrum is displayed.
+The optimized responses show a more uniform behavior.
+While resonances are inherent to the FDN structure, their prominence has been reduced and more evenly distributed.
+The effect of these changes on perceived coloration is further evaluated through a subjective test.
+Figure 6: Distribution of the modal excitation of an FDN with sizes $N=4,6,\,\text{and}\,\,8$ at the beginning (Init) and at the end of optimization (Optim), which has led to a decrease of the loudest modal excitation and to a narrowing of the distribution.
+Figure 7: Section of the magnitude response of the FDNs in Figure 6.
+The responses at initialization (Init) and at the end of optimization (Optim) have been shifted apart by 50 dB for better clarity.
+During optimization, the resonances have been reduced and more evenly distributed.
+Table 4: Standard deviation of the modal excitation before (Init) and after optimization (Optim) of different frameworks, each employing an FDN of size N=4 and identical set of delays.
 FDN type Init Optim
 DAFx23 7.8346 4.4802
-DiffFDN 7.8346 4.4518
-DiffFDN (HH) 8.5641 7.8765
-DiffFDN (SCAT) 7.6200 5.8822
+DiffFDN-O 7.8346 4.4518
+DiffFDN-HH 8.5641 7.8765
+DiffFDN-SCAT 7.6200 5.8822
 […]
@@
 6.1 Listening test procedure
-Figure 6: Listening test results on the analysized reverberation conditions: (a) lossless, (b) homogeneous decay, and (c) frequency-dependent decay.
+Figure 8: Listening test results on the analyzed reverberation conditions: (a) lossless, (b) homogeneous decay, and (c) frequency-dependent decay.
+The three DiffFDN conditions correspond to FDN configurations trained using the proposed optimization framework.
 The test followed the Multiple Stimuli with Hidden Reference and Anchor (MUSHRA) recommendation [33], and it was carried out using the web audio API-based experiment software webMUSHRA developed by International Audio Laboratories Erlangen [34].
@@
 The IRs of the FDNs were extracted from the mixing time, thereby eliminating the initial echo buildup from the evaluation.
-The section testing homogeneous decay, focused on frequency-independenty decaying IRs, with a $T_{60}$ value set at 2 seconds for all frequencies.
+The section testing homogeneous decay, focused on frequency-independently decaying IRs, with a $T_{60}$ value set at 2 s for all frequencies.
 A corresponding exponential decaying envelope with the same $T_{60}$ value was applied to the reference white Gaussian noise.
@@
 The frequency-dependent attenuation filter was designed following the steps outlined in Section 4.1.
-The test evaluated the coloration of the standard differentiable FDN (DiffFDN) proposed in this paper, the DiffFDN with scattering matrix (SCAT), and that with optimizable Householder matrix (HH).
+The test evaluated the coloration of the standard differentiable FDN with orthogonal feedback matrix (DiffFDN-O) proposed in this paper, the DiffFDN with scattering matrix (DiffFDN-SCAT), and that with optimizable Householder matrix (DiffFDN-HH).
 Each configuration was tested on a separate page where the number and lengths of the delays were constant, and only the feedback matrix and the input and output gains were altered.
 The FDN implementation of the Schroeder series allpass reverberator (SH) was the anchor, whereas the random orthogonal feedback matrix (RO) acted as a baseline.
-For the RO condition, the initial values of optimization of the DiffFDN were used.
+For the RO condition, the initial values of optimization of DiffFDN-O were used.
 The direct gain $d$ was set to zero in all cases.
+We excluded the DiffFDN trained with the framework presented in [9] since it would constitute a condition perceptually nearly identical to DiffFDN-O, and as such might skew the results.
+The primary contribution of this work is focused on improving training efficiency.
 In total, there were 18 listening test pages with six stimuli each.
@@
 The $p$-values for most paired conditions indicate significant differences among all pairs of results.
-However, there are a few exceptions observed in the distribution of DiffFDN and HH.
+However, there are a few exceptions observed in the distribution of DiffFDN-O and DiffFDN-HH.
 In these cases, the $p$-values exceed the alpha level across all conditions with decaying modes, suggesting a convergence in the scores assigned to the two conditions.
-Moreover, DiffFDN and HH reject the null hypothesis when paired with SCAT in the lossy cases for $N=8$, suggesting that for larger FDN sizes, the performance improvement provided by SCAT diminishes.
+Moreover, DiffFDN-O and DiffFDN-HH reject the null hypothesis when paired with DiffFDN-SCAT in the lossy cases for $N=8$, suggesting that for larger FDN sizes, the performance improvement provided by DiffFDN-SCAT diminishes.
 The IRs of the optimized FDNs always show a median greater than the initial condition (RO), proving the effectiveness of the optimization.
 In the late reverb analysis, Fig.
-6 (a), the DiffFDN has median ratings of 74, 82.5, and 70, respectively, for increasing FDN size.
-Only for $N=$ 8, the FDN with scattering matrix received ratings higher than DiffFDN with a median of 76.
-Introducing early reflections provides an advantage for the FDN with a scattering matrix (SCAT) compared to other conditions.
+8 (a), the DiffFDN-O has median ratings of 74, 82.5, and 70, respectively, for increasing FDN size.
+Only for $N=$ 8, the FDN with scattering matrix received ratings higher than DiffFDN-O with a median of 76.
+In the homogeneous decay and frequency-dependent decay tests the full IRs are taken into account.
+The scores given to the DiffFDN-SCAT condition suggest that the scattering matrix contributes to the smoothness of the early reflections as well as the tail.
 In general, spectral flatness is associated with temporal smoothness.
@@
 The confidence intervals for the results of all tests are relatively large.
-Although the objective results in Table 4 indicate a limited reduction in modal excitation standard deviation, the optimized Householder matrix was rated relatively close to the DiffFDN in the listening test.
+Although the objective results in Table 4 indicate a limited reduction in modal excitation standard deviation, the optimized Householder matrix was rated relatively close to the DiffFDN-O in the listening test.
 The distribution of modal excitation is one quantity that can aid in deducing the coloration of the reverberator.
@@
 In objective evaluations, we demonstrated that the proposed method reduces the width of the modal excitation distribution, thereby decreasing the prominence of the loudest modes.
-Listening test results further confirmed the success of this approach in attenuating artifacts commonly associated with undesired coloration, making the DiffFDN close to an ideally colorless, smooth reverberator.
+Listening test results further confirmed the success of this approach in attenuating artifacts commonly associated with undesired coloration, making the DiffFDN close to an ideally colorless, smooth late reverberation.
 Declarations
 Availability of data and materials
-The PyTorch implementation of the proposed method can be found on the online repository  [^1] https://github.com/gdalsanto/diff-fdn-colorless/.
+The PyTorch implementation of the proposed method, together with the code used to generate the figures, can be found on the online repository  [^1] https://github.com/gdalsanto/diff-fdn-colorless/.
 Listening test stimuli examples can be found on the dedicated website  [^2] http://research.spa.aalto.fi/publications/papers/eurasip-colorless-fdn/.